OpenStudy (anonymous):

Find the values of constants c and d that make the function continuous everywhere.

OpenStudy (anonymous):

F(x)= \[4x;x<-2\]\[cx+d;-2 \le x < 3\]\[-6x;x \ge 3\]

OpenStudy (anonymous):

G(x)=\[3x;x<2\]\[cx^2+d;2 \le x <4\]\[5x-2;x \ge 4\]

OpenStudy (anonymous):

replace \(x\) by \(2\) in the first two expressions and get \(6=4c+d\)

OpenStudy (anonymous):

then replace \(x\) by \(4\) in the second two and get \(16c+d=18\)

OpenStudy (anonymous):

solve the system for \(c\) and \(d\)

OpenStudy (anonymous):

Thank you Satellite for the help.

OpenStudy (anonymous):

yw